OpenStudy (anonymous):

0.12 * 10 -9 as a decimal

OpenStudy (anonymous):

-7.8

OpenStudy (anonymous):

I mean 0.12 * 10^-9

jhonyy9 (jhonyy9):

0,00000000012

OpenStudy (anonymous):

Order of operations says you multiply before subtracting = (0.12 * 10) - 9 =(1.2) -9 = -7.8

OpenStudy (anonymous):

just read your mis type message...

OpenStudy (anonymous):

Scientific notation....take the decimal and move to the LEFT 9 places .12 to the left 9 places = .00000000012

OpenStudy (anonymous):

if it was a POSITIVE EXPONET.... you would move the decimal to the RIGHT 9 places. get it?